OpenStudy (samigupta8):

A straight line intersects the same branch of general hyperbola in P1 and P2 and meets its asymptotes in Q1 and Q2 .Then P1Q2-P2Q1 is equal to....

OpenStudy (samigupta8):

@hartnn @parthkohli pls...do help..

Parth (parthkohli):

just because of the nature of the information provided in the problem, i would say the answer is zero. lol

OpenStudy (samigupta8):

How?

Parth (parthkohli):

Because of symmetry it's obvious for any vertical line. But let's solve it in a proper manner. Asymptotes:\[x^2/a^2 - y^2/b^2 = 0\]Hyperbola\[x^2/a^2 - y^2/b^2 = 1\]

OpenStudy (samigupta8):

Then....

Parth (parthkohli):

Then take \(y= mx+c\) and solve.

Parth (parthkohli):

Not recommended though

Parth (parthkohli):

But it's really obvious for vertical lines because of symmetry about axes.

OpenStudy (samigupta8):

What about any general line ??

Parth (parthkohli):

sorry i'll be away for a while

OpenStudy (samigupta8):

It's alryt....

OpenStudy (hauntedwoodsgal):

Hi still need help or are you being helped??

OpenStudy (samigupta8):

No, i need help....

OpenStudy (hauntedwoodsgal):

I have to agree with @ParthKholi

OpenStudy (samigupta8):

Can't we do it assuming general line

OpenStudy (hauntedwoodsgal):

The answer is zero I got it also

OpenStudy (samigupta8):

U also assumed vertical line only....

OpenStudy (hauntedwoodsgal):

yes

OpenStudy (samigupta8):

Bt y not other line

OpenStudy (hauntedwoodsgal):

Brb

OpenStudy (samigupta8):

Ok...

OpenStudy (hauntedwoodsgal):

Ok....Back

OpenStudy (samigupta8):

Yaa so now would you tell me why not in generalized way?

OpenStudy (hauntedwoodsgal):

Oh that's hard and im not used to doing it in a not generalized way but I will try.......

OpenStudy (samigupta8):

Yup...surely....

OpenStudy (hauntedwoodsgal):

hmmmmmm

OpenStudy (hauntedwoodsgal):